mrs lopez wants to tile your kitchen floor. her room is 10 ft by 12 ft. how many 1 foot square tile does she need to cover her floor.
i got that she will need 120 tiles since the area of the room is 120 i am not sure with this answer
pasrt b says that mrs lopez changes her mindand choosesz smaller tiles that are only 6 inches on each side. how many tiles does she need

Answer by mananth(16946)

120 tiles is right
..
part II
area of tile = 1/2 * 1/2 = 1/4
total area of floor = 120
120/(1/4)
480 tiles